My word problem is: Sandy's electric bill is always 5 times as large as her gas bill. If the total of the two bills is $174, how much was Sandy's electric bill?
I was thinking 5x=174 and divide it and got 34.8

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

My word problem is: Sandy's electric bill is always 5 times as large as her gas bill. If the total of the two bills is $174, how much was Sandy's electric bill?
I was thinking 5x=174 and divide it and got 34.8
let the gas bill be x
electric bill will be 5x
5x+x=174
6x=174
x=174/6
= $29 which is the gas bill
The elctricity bill will be 5times x
ie: 5*29= 145

Answer by Alan3354(69443)

E + G = 174
E = 5G